# Privacy-aware Berrut Approximated Coded Computing for Federated Learning

### v1

@@ IV PBSS: PBACC with Secret Sharing
While PBACC provides quantifiable privacy for distributed computations, it is not yet valid for FL, since it only supports a single data owner (the master). The workers are just delegated nodes for the computation tasks. In contrast, in FL the roles are reversed: the clients perform local training and are also the data owners, whereas the aggregator node is the central entity doing the model aggregation. Therefore, to be useful for general FL frameworks, PBACC needs to be enhanced to a multi-input secret sharing configuration, one that allows a set of nodes to collaboratively compute a function, either training or aggregation, whose final results can be decoded by the central aggregator. After the protocol is defined, it is also essential to measure its communication and computation costs in each phase, so that it proves scalable in addition to privacy-preserving. In the following, the Private BACC Secret Sharing (PBSS) protocol is presented (Section IV-A), and its computational complexity is analyzed (Section IV-B).
@@ IV-A Private BACC Secret Sharing (PBSS) protocol
Suppose a collection of $N$ nodes, each one having a private data vector $\mathbf{X}^{(i)}=\begin{pmatrix}X_{0}^{(i)}&\ldots&X_{K-1}^{(i)}\end{pmatrix}^{T}$, where the input data $\mathbf{X}^{(i)}\in\mathbb{R}^{K\times L}$ is owned by node $i$, for $i\in\{0,\dots,N-1\}$. The PBSS consists of three phases, that are detiled as follows:
@@
$v_{\mathbf{X}^{(i)}}(z)$ $=\sum_{j=0}^{K-1}\frac{\frac{(-1)^{j}}{(z-\alpha_{j})}}{\sum_{k=0}^{K+T-1}\frac{(-1)^{k}}{(z-\alpha_{k})}}X_{j}^{(i)}$ (5)
$+\sum_{j=0}^{T-1}\frac{\frac{(-1)^{j+K}}{(z-\alpha_{j})}}{\sum_{k=0}^{K+T-1}\frac{(-1)^{j}}{(z-\alpha_{k})}}R_{j}^{(i)},$
for some distinct interpolation points associated to the the data $\mathbf{X}^{(i)}$, $\boldsymbol{\alpha}=(\alpha_{0},\dots,\alpha_{K-1})\in\mathbb{R}^{K}$, which we choose again as the Chebyshev points of first kind $\alpha_{j}=\cos\left(\frac{(2j+1)\pi}{2K}\right)$, and some distinct interpolation points associated to the $R^{(i)}_{j}$, $\alpha_{K},\dots,\alpha_{K+T-1}\in\mathbb{R}$, which are selected as the shifted Chebyshev points of first kind $\alpha_{K+j}=b+\cos\left(\frac{(2j+1)\pi}{2K}\right)$. The rational function $v_{\mathbf{X}^{(i)}}(z)$ is evaluated at the set $\{z_{j}\}$, for $j=0,\dots,N-1$, using $z_{j}=\cos(\frac{j\pi}{N-1})$, the Chebyshev points of second kind. Finally, the random coefficients $R_{j}^{(i)}$ are distinct $1\times L$ vectors drawn from a Gaussian distribution $\mathcal{N}(0,\frac{\sigma_{n}^{2}}{T})$. The evaluation $v_{\mathbf{X}^{(i)}}(z_{j})$ is the share created from the client node $i$ and sent to the node $j$, as explained below. The sharing phase is represented in Fig. 2a assuming there are three workers owning their own private data set ($X_{j}$) and locally computing a local AI model. As it is shown, the aggregator sends the global model ($\theta$) to the nodes, so they can complete a new round of their learning task. The parameters of the global model are codified, using the function $u()$ established in eq. (5), and shared with the three nodes. After this codification part, the PBSS scheme acts: each node $j$ receives the shares created using the evaluation point $z_{j}$, and computes the aggregation function $f()$ over the codified parameters $u_{\theta^{\prime}_{i}}(z_{j})$. In this case, the input data to be secured for the node $i$ is $\theta^{\prime}_{i}$, the model updates after the local computation is done.
Phase 2 - Computation and communication: In this phase, the node $i$ calculates an arbitrary function $f$ using a set of polynomial quotient evaluations shared from the rest of nodes $\{v_{\mathbf{X}^{(0)}}(z_{i}),u_{\mathbf{X}^{(1)}}(z_{i}),\dots,u_{\mathbf{X}^{(N-1)}}(z_{i})\}$, where $u_{\mathbf{X}^{(j)}}(z_{i})$ is the evaluation of the rational function corresponding to the input $X^{(j)}$ owned by the $j$-th node, and shared with node $i$. The client node $i$ computes $f\bigl{(}v_{\mathbf{X}^{(j)}}(z_{i})\bigr{)},\quad j=\{0,\dots,N-1\}$ and sends the result to the master node. This phase is schematically represented in Fig. 2b, where the obtained results $f(u_{\theta^{\prime}_{i}}(z_{j}))$ are sent to the aggregator.
@@
$r_{\mathrm{Berrut},\mathcal{F}}(z)=\sum_{i=0}^{n}\frac{\frac{(-1)^{i}}{(z-\tilde{z}_{i})}}{\sum_{j=0}^{n}\frac{(-1)^{j}}{(z-\tilde{z}_{j})}}f\bigl{(}v(\tilde{z}_{i})\bigr{)},$ (6)
where $\tilde{z}_{i}\in\mathcal{S}=\{\cos\frac{j\pi}{N-1},j\in\mathcal{F}\}$ are the interpolation points. At this moment, the master node finds the approximation $f(X^{(i)}_{j})\approx r_{\mathrm{Berrut},\mathcal{F}}(\alpha_{j})$, for all $j\in\{0,\dots,K-1\}$ and $i\in\{0,\dots,N-1\}$.
Fig. 2c shows the reconstruction phase for a secure aggregation. When the aggregator receives the results $f(u_{\theta^{\prime}_{0}}(z_{j}),$ $u_{\theta^{\prime}_{1}}(z_{j}),$ $\dots$,… , $u_{\theta^{\prime}_{N-1}}(z_{j}))$ from each node, it can apply the decoding function $r()$ to decode the global model. After that, the aggregator would compute some arbitrary aggregation algorithm, e.g., SGD or any of its many variants in the literature (FedAvg [42], SCAFFOLD [43] or FedGen [44]). Finally, the aggregator sends the updated global model to the worker nodes, so the process can be repeated until convergence is achieved.
@@ IV-B2 Computation and Communication
In this phase, client node $i$ computes $f(u(z_{i}))$, for the arbitrary function $f$, with $u(z_{i})$ a coded share of size $1\times L$, thus with a $\mathcal{O}(L)$ complexity. This is followed by transmission of the result wto the master node, so $N$ messages are exchanged at most, each one with size proportional to $L$.
@@ IV-B3 Reconstruction:
Since $\text{len}(Z)=K+T$, the total computational complexity is $\mathcal{O}(K(2n+L)+3K+3T+N+L)$. For the communication complexity, $N(N-1)+N$ messages of a size proportional to $L$ are exchanged at most.
V Matrix multiplication with BSS
@@
However, our method has two main limitations: (i) the inability to reconstruct the final result when the number of worker nodes is lower than the number of rows of the input matrices, and (ii) high communication costs in comparison to BSS. As massive input matrices are common in machine learning, these two limitations might hinder the utility of the technique for large models. We mitigate the aforementioned problems using two ideas: first, by encoding multiple rows of a matrix in the same interpolation point; and secondly, by using block matrix multiplication for large matrices, where the input matrix in partitioned in vertical blocks.
@@ V-A Approximate Matrix Multiplication without privacy
The input matrices $A$ and $B$ are encoded using the same rational function $u_{i}(z)=q_{i}(z)X_{i}$, with $X_{1}=A$, $X_{2}=B$ and
$q_{i}(z)=\frac{\frac{(-1)^{i}}{(z-\alpha_{i})}}{\sum_{j=0}^{K-1}\frac{(-1)^{j}}{(z-\alpha_{j})}},\quad i=1,2.$
@@
Node $k$ receives two coded shares $u^{(A)}(z_{k})=u_{1}(z_{k})$ and $u^{(B)}(z_{k})=u_{2}(z_{k})$ corresponding to the two input matrices $A$ and $B$ $A=\begin{pmatrix}A_{0}&A_{1}&\ldots&A_{K-1}\end{pmatrix}^{T}$, $B=\begin{pmatrix}B_{0}&B_{1}&\ldots&B_{K-1}\end{pmatrix}^{T}$ for $A_{i}\in\mathbb{R}^{1\times d}$, $B_{i}\in\mathbb{R}^{\times l\ell}$. Once the node $k$ has the operands, it computes the matrix multiplication so $f(u^{(A)}(z_{k}),u^{(B)}(z_{k}))=u^{(A)}(z_{k})\bigl{(}u^{(B)}(z_{k})\bigr{)}^{T}$.
@@
Node $k$ divides each column $j$ of $u^{(A)}(z_{k})\bigl{(}u^{(B)}(z_{k})\bigr{)}^{T}$ by the factor $q_{j}(z_{k})$, so
$g(u(z_{k}))=\begin{pmatrix}\frac{u_{0}^{(A)}(z_{k})\bigl{(}u_{0}^{(B)}(z_{k})\bigr{)}^{T}}{q_{0}(z_{k})}&\ldots&\frac{u_{0}^{(A)}(z_{k})\bigl{(}u_{K-1}^{(B)}(z_{k})\bigr{)}^{T}}{q_{K-1}(z_{k})}\\ \frac{u_{1}^{(A)}(z_{k})\bigl{(}u_{0}^{(B)}(z_{k})\bigr{)}^{T}}{q_{0}(z_{k})}&\ldots&\frac{u_{1}^{(A)}(z_{k})\bigl{(}u_{K-1}^{(B)}(z_{k})\bigr{)}^{T}}{q_{K-1}(z_{k})}\\ \vdots&\ddots&\vdots\\ \frac{u_{K-1}^{(A)}(z_{k})\bigl{(}u_{0}^{(B)}(z_{k})\bigr{)}^{T}}{q_{0}(z_{k})}&\ldots&\frac{u_{K-1}^{(A)}(z_{k})\bigl{(}u_{K-1}^{(B)}(z_{k})\bigr{)}^{T}}{q_{K-1}(z_{k})}\\ \end{pmatrix}$ (7)
for all $j=0,\dots,K-1$. Since $u_{i}^{(A)}(z_{k})=q_{i}(z_{k})A_{i}$, we can rewrite (7) as
$g(u(z_{k}))=\begin{pmatrix}\frac{q_{0}(z_{k})A_{0}q_{0}(z_{k})B_{0}^{T}}{q_{0}(z_{k})}&\ldots&\frac{q_{0}(z_{k})A_{0}q_{K-1}(z_{k})B_{K-1}^{T}}{q_{K-1}(z_{k})}\\ \frac{q_{1}(z_{k})A_{1}q_{0}(z_{k})B_{0}^{T}}{q_{0}(z_{k})}&\ldots&\frac{q_{1}(z_{k})A_{1}q_{K-1}(z_{k})B_{K-1}^{T}}{q_{K-1}(z_{k})}\\ \vdots&\ddots&\vdots\\ \frac{q_{K-1}(z_{j})A_{K-1}q_{0}(z_{k})B_{0}^{T}}{q_{0}(z_{k})}&\ldots&\frac{q_{K-1}(z_{k})A_{K-1}q_{K-1}(z_{k})B_{K-1}^{T}}{q_{K-1}(z_{k})}\end{pmatrix},$
@@
$\begin{pmatrix}q_{0}(z_{k})A_{0}B_{0}^{T}&\ldots&q_{0}(z_{k})A_{0}B_{K-1}^{T}\\ q_{1}(z_{k})A_{1}B_{0}^{T}&\ldots&q_{1}(z_{k})A_{1}B_{K-1}^{T}\\ \vdots&\ddots&\vdots\\ q_{K-1}(z_{k})A_{K-1}B_{0}^{T}&\ldots&q_{K-1}(z_{k})A_{K-1}B_{K-1}^{T}\end{pmatrix}=$
$\operatorname{diag}\bigl{(}q_{0}(z_{k}),\dots,q_{K-1}(z_{k})\bigr{)}\begin{pmatrix}A_{0}B_{0}^{T}&\ldots&A_{0}B_{K-1}^{T}\\ A_{1}B_{0}^{T}&\ldots&A_{1}B_{K-1}^{T}\\ \vdots&\ddots&\vdots\\ A_{K-1}B_{0}^{T}&\ldots&A_{K-1}B_{K-1}^{T}\end{pmatrix}$
@@
$f(u(z_{k}))=\begin{pmatrix}\sum_{i=0}^{K-1}q_{i}(z_{k})A_{i}B_{0}^{T}&\ldots&\sum_{i=0}^{K-1}q_{i}(z_{k})A_{i}B_{K-1}^{T}\end{pmatrix},$ (8)
@@
Recalling the explanation of BACC in Section II, what the decoding operation does is to approximately calculate the value of the function at the points $\alpha_{0},\alpha_{1},\dots,\alpha_{K-1}$, so $r_{\mathrm{Berrut},\mathcal{F}}(\alpha_{i})\approx f(X_{i})$. In order to achieve successful decoding work, it is required that $u(\alpha_{i})=X_{i}$. As the shares have the form $\sum_{i=0}^{K-1}q_{i}(z_{j})X_{i}$, this is equivalent to saying that, when $u$ is evaluated in $\alpha_{i}$, $q_{i}(z_{j})=1$, and the rest of $q_{k}(z_{j})=0$ for all $k\neq i$. For applying this to our matrix multiplication, we can easily check using (8) that if $q_{i}(z_{j})=1$ and $q_{k}(z_{j})=0$ for all $k\neq i$, the resulting row is $f\bigl{(}u(z_{j})\bigr{)}[\alpha_{i}]=\begin{pmatrix}A_{i}B_{0}^{T}&A_{i}B_{1}^{T}&\ldots&A_{i}B_{K-1}^{T}\end{pmatrix}$, which is exactly the value of row $i$ of $AB^{T}$, and this holds for any $i$.

### v2

@@ IV PBSS: PBACC with Secret Sharing
While PBACC provides quantifiable privacy for distributed computations, it is not yet valid for FL, since it only supports a single data owner (the master). The workers are just delegated nodes for the computation tasks. In contrast, in FL the roles are reversed: the clients perform local training and are also the data owners, whereas the aggregator node is the central entity doing the model aggregation. Therefore, to be useful for general FL frameworks, PBACC needs to be enhanced to a multi-input secret sharing configuration, one that allows a set of nodes to collaboratively compute a function, either training or aggregation, whose final results can be decoded by the central aggregator. After the protocol is defined, it is also essential to measure its communication and computation costs in each phase, so that it proves to be scalable in addition to privacy-preserving. In the following, the Private BACC Secret Sharing (PBSS) protocol is presented (Section IV-A), and its computational complexity is analyzed (Section IV-B). Finally, we have added a comparison of the proposed scheme with some SMPC and HE approaches for non-linear function computations (IV-C).
@@ IV-A Private BACC Secret Sharing (PBSS) protocol
Suppose a collection of $N$ nodes, each one having a private data vector $\mathbf{X}^{(i)}=\begin{pmatrix}X_{0}^{(i)}&\ldots&X_{K-1}^{(i)}\end{pmatrix}^{T}$, where the input data $\mathbf{X}^{(i)}\in\mathbb{R}^{K\times L}$ is owned by node $i$, for $i\in\{0,\dots,N-1\}$. The PBSS consists of three phases, that are detailed as follows:
@@
$u_{\mathbf{X}^{(i)}}(z)$ $=\sum_{j=0}^{K-1}\frac{\frac{(-1)^{j}}{(z-\alpha_{j})}}{\sum_{k=0}^{K+T-1}\frac{(-1)^{k}}{(z-\alpha_{k})}}X_{j}^{(i)}+\sum_{j=0}^{T-1}\frac{\frac{(-1)^{j+K}}{(z-\alpha_{j})}}{\sum_{k=0}^{K+T-1}\frac{(-1)^{j}}{(z-\alpha_{k})}}R_{j}^{(i)},$ (5)
for some distinct interpolation points associated to the the data $\mathbf{X}^{(i)}$, $\boldsymbol{\alpha}=(\alpha_{0},\dots,\alpha_{K-1})\in\mathbb{R}^{K}$, which we choose again as the Chebyshev points of first kind $\alpha_{j}=\cos\left(\frac{(2j+1)\pi}{2K}\right)$, and some distinct interpolation points associated to the $R^{(i)}_{j}$, $\alpha_{K},\dots,\alpha_{K+T-1}\in\mathbb{R}$, which are selected as the shifted Chebyshev points of first kind $\alpha_{K+j}=b+\cos\left(\frac{(2j+1)\pi}{2K}\right)$. The rational function $u_{\mathbf{X}^{(i)}}(z)$ is evaluated at the set $\{z_{j}\}$, for $j=0,\dots,N-1$, using $z_{j}=\cos(\frac{j\pi}{N-1})$, the Chebyshev points of second kind. Finally, the random coefficients $R_{j}^{(i)}$ are distinct $1\times L$ vectors drawn from a Gaussian distribution $\mathcal{N}(0,\frac{\sigma_{n}^{2}}{T})$. The evaluation $u_{\mathbf{X}^{(i)}}(z_{j})$ is the share created from the client node $i$ and sent to the node $j$, as explained below. The sharing phase is represented in Fig. 3a assuming there are three workers owning their own private data set ($X_{j}$) and locally computing an AI model. As it is shown, the aggregator sends the global model ($\theta$) to the nodes, so they can complete a new round of their learning task. The parameters of the global model are codified, using the function $u()$ established in eq. (5), and shared with the three nodes. After this codification part, the PBSS scheme acts: each node $j$ receives the shares created using the evaluation point $z_{j}$, and computes the aggregation function $f()$ over the codified parameters $u_{\theta^{\prime}_{i}}(z_{j})$. In this case, the input data to be secured for the node $i$ is $\theta^{\prime}_{i}$, the model updates after the local computation is done.
Phase 2 - Computation and communication: In this phase, the node $i$ calculates an arbitrary function $f$ using a set of polynomial quotient evaluations shared from the rest of nodes $\{u_{\mathbf{X}^{(0)}}(z_{i}),u_{\mathbf{X}^{(1)}}(z_{i}),\dots,u_{\mathbf{X}^{(N-1)}}(z_{i})\}$, where $u_{\mathbf{X}^{(j)}}(z_{i})$ is the evaluation of the rational function corresponding to the input $X^{(j)}$ owned by the $j$-th node, and shared with node $i$. The client node $i$ computes $f\bigl{(}u_{\mathbf{X}^{(0)}}(z_{i}),u_{\mathbf{X}^{(1)}}(z_{i}),\dots,u_{\mathbf{X}^{(N-1)}}(z_{i})\bigr{)}$ and sends the result to the master node. This phase is schematically represented in Fig. 3b, where the obtained results $f\bigl{(}u_{\theta^{\prime}_{0}}(z_{j}),u_{\theta^{\prime}_{1}}(z_{j}),u_{\theta^{\prime}_{2}}(z_{j})\bigr{)}$ are sent to the aggregator.
@@
$r_{\mathrm{Berrut},\mathcal{F}}(z)=\sum_{i=0}^{n}\frac{\frac{(-1)^{i}}{(z-\tilde{z}_{i})}}{\sum_{j=0}^{n}\frac{(-1)^{j}}{(z-\tilde{z}_{j})}}f\bigl{(}u_{\mathbf{X}^{(0)}}(\tilde{z}_{i}),u_{\mathbf{X}^{(1)}}(\tilde{z}_{i}),\dots,u_{\mathbf{X}^{(N-1)}}(\tilde{z}_{i})\bigr{)},$ (6)
where $\tilde{z}_{i}\in\mathcal{S}=\{\cos\frac{j\pi}{N-1},j\in\mathcal{F}\}$ are the evaluation points of the fastest nodes. At this moment, the master node finds the approximation $f\bigl{(}X^{(0)}_{j},X^{(1)}_{j},\dots,X^{(N-1)}_{j}\bigr{)}\approx r_{\mathrm{Berrut},\mathcal{F}}(\alpha_{j})$, for all $j\in\{0,\dots,K-1\}$.
Fig. 3c shows the reconstruction phase for a secure aggregation. When the aggregator receives the results $f(u_{\theta^{\prime}_{0}}(z_{j}),u_{\theta^{\prime}_{1}}(z_{j}),u_{\theta^{\prime}_{2}}(z_{j}))$ from each node, it can apply the decoding function $r()$ to decode the global model. After that, the aggregator would compute some arbitrary aggregation algorithm, e.g., SGD or any of its many variants in the literature (FedAvg [43], SCAFFOLD [44] or FedGen [45]). Finally, the aggregator sends the updated global model to the worker nodes, so the process can be repeated until convergence is achieved.
@@ IV-B2 Computation and Communication
In this phase, client node $i$ computes $f\bigl{(}u_{X^{(0)}}(z_{i}),u_{X^{(1)}}(z_{i}),\dots,u_{X^{(N-1)}}(z_{i})\bigr{)}$, for the arbitrary function $f$, with $u_{X^{(j)}}(z_{i})$ a coded share of size $1\times L$ received by node $j$, thus with a $\mathcal{O}(NL)$ complexity. This is followed by transmission of the result to the master node, so $N$ messages are exchanged at most, each one with size proportional to $L$.
@@ IV-B3 Reconstruction:
Since $\text{len}(Z)=K+T$, the total computational complexity is $\mathcal{O}(K(2n+L)+3K+3T+N+NL)$. For the communication complexity, $N(N-1)+N$ messages of a size proportional to $L$ are exchanged at most.
IV-C Comparison with HE and SMPC schemes
TABLE I: Secure non-linear distributed computing schemes comparison.
protocol privacy precision functions
cons_ltz [46] Information-theoretic Exact ReLU
lss_ltz [46] Information-theoretic Exact ReLU
[47] for Softmax Information-theoretic Exact Softmax
[47] for Sigmoid Information-theoretic Exact Sigmoid
[48] Computational Approximate ReLU
PBSS Bounded Leakage Approximate Any
In [46], the authors propose some protocols based on improvements of Rabbit [49] for the computation of ReLUs, and test them for different configurations. Fastest protocols cons_ltz and lsss_ltz require several rounds of communication, $R=4$ and $R=7$ respectively, with $N^{2}$ messages exchanged for each of them. PBSS only requires $N^{2}$ messages at most with reduced size ($1\times L$) for any non-linear function computation. Regarding computation complexity, PBSS only needs to compute the ReLU itself for reduced inputs of size $1\times L$, while these protocols have to evaluate a Garbled Circuit of approximately $G=200$ gates over normal size inputs, which makes PBSS also outperform it. It is also worth mentioning that some of these protocols require circuit optimizers to reduce the number of communication rounds, which impose a really high memory usage to load the dependency graph. In [47], they propose a SMPC framework for the efficient computation of Softmax and Sigmoid. For the case of Softmax, $2r$ rounds are required, being $r$ the number of iterations (increasing $r$ results in greater accuracy, e.g., $r=\{8,16,32,64\}$), with a complexity of $(8mbr+2b)rKL$, being $m$ the number of classes, and $b$ the bit-length of fixed point numbers. For the case of Sigmoid, only $1$ round is required with a complexity of $12bKL$ at most. Similarly to what happens with Softmax, PBSS outperforms them in communication complexity due to only requiring one round, and reducing the size of the messages. Regarding computation, since we can directly compute the non-linear functions without making any approximation for a polynomial, PBSS also outperforms them. In [48], they propose a very efficient approximation of the ReLU function for homomorphic evaluation that requires from $P\in[7,8,\dots,30]$ multiplications depending on the precision required. Since they are testing it with CKKS, the size of the ciphertext scales linearly with the multiplicative depth, which makes PBSS outperforms it in computation times. Regarding communication, this solution will always outperform PBSS, since it does not require to perform several rounds of communication for evaluation or decryption. Additionally, it is important to highlight that we are comparing our solution with schemes specifically dedicated and optimized for the execution of one or two types of non-linear functions, while PBSS can compute any arbitrary function. The main downsides of PBSS against all solutions compared is that:
  1. 1.
    It provides a weaker notion of privacy (bounded privacy leakage guarantee), contrary to the other solutions, that provide either information-theoretic privacy or computational privacy guarantees.
  2. 2.
    It presents a high precision loss in small networks that require to share massive inputs. Regarding this issue, we have proposed two features to tackle it in V-D.
V Matrix multiplication with PBSS
@@
However, our method has two main limitations: (i) the error of the reconstruction is too high when the number of worker nodes is lower than the number of rows of the input matrices (this is inherited from the original BACC scheme), and (ii) high communication costs in comparison to PBSS without privacy, from now on, BSS. As massive input matrices are common in machine learning, these two limitations might hinder the utility of the technique for large models. We mitigate the aforementioned problems using two ideas: first, by encoding multiple rows of a matrix in the same interpolation point; and secondly, by using block matrix multiplication for large matrices, where the input matrix is partitioned in vertical blocks.
@@ V-A Approximate Matrix Multiplication without privacy
Each of the rows of the input matrices $A$ and $B$ are encoded using the rational function $u_{i}(z)=\hat{q}_{i}(z)X_{i}$, being $X_{i}$ the row $i$ of the matrix to encode, and
$\hat{q}_{i}(z)=\frac{\frac{(-1)^{i}}{(z-\alpha_{i})}}{\sum_{j=0}^{K-1}\frac{(-1)^{j}}{(z-\alpha_{j})}},\quad i=0,1,\dots,K-1.$
@@
Node $k$ receives two coded shares $u^{(A)}(z_{k})$ and $u^{(B)}(z_{k})$ corresponding to the two input matrices $A=\begin{pmatrix}A_{0}&A_{1}&\ldots&A_{K-1}\end{pmatrix}^{T}$ and $B=\begin{pmatrix}B_{0}&B_{1}&\ldots&B_{K-1}\end{pmatrix}^{T}$ for $A_{i}\in\mathbb{R}^{1\times d}$, $B_{i}\in\mathbb{R}^{1\times\ell}$. Once the node $k$ has the operands, it computes the matrix multiplication so $f(u^{(A)}(z_{k}),u^{(B)}(z_{k}))=u^{(A)}(z_{k})\bigl{(}u^{(B)}(z_{k})\bigr{)}^{T}$.
@@
Node $k$ divides each column $j$ of $u^{(A)}(z_{k})\bigl{(}u^{(B)}(z_{k})\bigr{)}^{T}$ by the factor $\hat{q}_{j}(z_{k})$, so
$g(u(z_{k}))=\begin{pmatrix}\frac{u_{0}^{(A)}(z_{k})\bigl{(}u_{0}^{(B)}(z_{k})\bigr{)}^{T}}{\hat{q}_{0}(z_{k})}&\ldots&\frac{u_{0}^{(A)}(z_{k})\bigl{(}u_{K-1}^{(B)}(z_{k})\bigr{)}^{T}}{\hat{q}_{K-1}(z_{k})}\\ \frac{u_{1}^{(A)}(z_{k})\bigl{(}u_{0}^{(B)}(z_{k})\bigr{)}^{T}}{\hat{q}_{0}(z_{k})}&\ldots&\frac{u_{1}^{(A)}(z_{k})\bigl{(}u_{K-1}^{(B)}(z_{k})\bigr{)}^{T}}{\hat{q}_{K-1}(z_{k})}\\ \vdots&\ddots&\vdots\\ \frac{u_{K-1}^{(A)}(z_{k})\bigl{(}u_{0}^{(B)}(z_{k})\bigr{)}^{T}}{\hat{q}_{0}(z_{k})}&\ldots&\frac{u_{K-1}^{(A)}(z_{k})\bigl{(}u_{K-1}^{(B)}(z_{k})\bigr{)}^{T}}{\hat{q}_{K-1}(z_{k})}\\ \end{pmatrix}$ (7)
for all $j=0,\dots,K-1$. Since $u_{i}^{(A)}(z_{k})=\hat{q}_{i}(z_{k})A_{i}$, we can rewrite (7) as
$g(u(z_{k}))=\begin{pmatrix}\frac{\hat{q}_{0}(z_{k})A_{0}\hat{q}_{0}(z_{k})B_{0}^{T}}{\hat{q}_{0}(z_{k})}&\ldots&\frac{\hat{q}_{0}(z_{k})A_{0}\hat{q}_{K-1}(z_{k})B_{K-1}^{T}}{\hat{q}_{K-1}(z_{k})}\\ \frac{\hat{q}_{1}(z_{k})A_{1}\hat{q}_{0}(z_{k})B_{0}^{T}}{\hat{q}_{0}(z_{k})}&\ldots&\frac{\hat{q}_{1}(z_{k})A_{1}\hat{q}_{K-1}(z_{k})B_{K-1}^{T}}{\hat{q}_{K-1}(z_{k})}\\ \vdots&\ddots&\vdots\\ \frac{\hat{q}_{K-1}(z_{j})A_{K-1}\hat{q}_{0}(z_{k})B_{0}^{T}}{\hat{q}_{0}(z_{k})}&\ldots&\frac{\hat{q}_{K-1}(z_{k})A_{K-1}\hat{q}_{K-1}(z_{k})B_{K-1}^{T}}{\hat{q}_{K-1}(z_{k})}\end{pmatrix},$
@@
$\begin{pmatrix}\hat{q}_{0}(z_{k})A_{0}B_{0}^{T}&\ldots&\hat{q}_{0}(z_{k})A_{0}B_{K-1}^{T}\\ \hat{q}_{1}(z_{k})A_{1}B_{0}^{T}&\ldots&\hat{q}_{1}(z_{k})A_{1}B_{K-1}^{T}\\ \vdots&\ddots&\vdots\\ \hat{q}_{K-1}(z_{k})A_{K-1}B_{0}^{T}&\ldots&\hat{q}_{K-1}(z_{k})A_{K-1}B_{K-1}^{T}\end{pmatrix}=\operatorname{diag}\bigl{(}\hat{q}_{0}(z_{k}),\dots,\hat{q}_{K-1}(z_{k})\bigr{)}\begin{pmatrix}A_{0}B_{0}^{T}&\ldots&A_{0}B_{K-1}^{T}\\ A_{1}B_{0}^{T}&\ldots&A_{1}B_{K-1}^{T}\\ \vdots&\ddots&\vdots\\ A_{K-1}B_{0}^{T}&\ldots&A_{K-1}B_{K-1}^{T}\end{pmatrix}$
@@
$f(u(z_{k}))=\begin{pmatrix}\sum_{i=0}^{K-1}\hat{q}_{i}(z_{k})A_{i}B_{0}^{T}&\ldots&\sum_{i=0}^{K-1}\hat{q}_{i}(z_{k})A_{i}B_{K-1}^{T}\end{pmatrix},$ (8)
@@
Recalling the explanation of BACC in Section III, what the decoding operation does is to approximately calculate the value of the function at the points $\alpha_{0},\alpha_{1},\dots,\alpha_{K-1}$, so $r_{\mathrm{Berrut},\mathcal{F}}(\alpha_{i})\approx f(X_{i})$. In order to achieve successful decoding work, it is required that $u(\alpha_{i})=X_{i}$. As the shares have the form $\sum_{i=0}^{K-1}\hat{q}_{i}(z_{j})X_{i}$, this is equivalent to saying that, when $u$ is evaluated in $\alpha_{i}$, $\hat{q}_{i}(z_{j})=1$, and the rest of $\hat{q}_{k}(z_{j})=0$ for all $k\neq i$. For applying this to our matrix multiplication, we can easily check using (8) that if $\hat{q}_{i}(z_{j})=1$ and $\hat{q}_{k}(z_{j})=0$ for all $k\neq i$, the resulting row is $f\bigl{(}u(z_{j})\bigr{)}[\alpha_{i}]=\begin{pmatrix}A_{i}B_{0}^{T}&A_{i}B_{1}^{T}&\ldots&A_{i}B_{K-1}^{T}\end{pmatrix}$, which is exactly the value of row $i$ of $AB^{T}$, and this holds for any $i$.